# Examining the Prevalence and Dynamics of AI-Generated Media in Art Subreddits

### v1

@@ Introduction
Figure 2: The cumulative distribution of posts across different upvote ratio values (x-axis), and the corresponding cumulative percentage of posts (y-axis) for different kinds of post (left vs right) and communities (top vs bottom).
@@
Given this observation, we further focused on the dynamics around Suspected AI posts, posts where a non-author commenter claimed or implied that the image attached to the post was generated using AI. We looked at the speed with which these claims emerged and the content of these claims. We began by calculating the time elapsed from the publication of the post until its first AI-claiming comment. Our analysis revealed that approximately 20% of the Suspected AI posts had received their first AI-claiming comment within one hour from posting. However, to compare between posts in the different types of communities, we had to control for the popularity of the post or its likelihood to be seen. To do that, we used the time window between the first comment on each Suspected AI post, and the first AI-claiming comment on that post. This between-comments data is presented in Figure 3, which shows the time between the comments (X-axis), and the cumulative number of posts (Y-axis). For instance, the leftmost data point in the figure indicates that for both community types, approximately 20% of the AI-claiming comments were made as the first comments on the posts (the X-axis tick mark of “1 minute” denoting “within one minute”, and mostly captured the first comment). The figure shows that these comments appear more quickly in AI-disallowed communities. For example, around 40% of the AI-claiming comments in these communities (red line) were published within one hour of the first comment. In AI-neutral communities, this speed of action occurred for less than 30% of the posts. The figure also shows that after approximately 36 hours, the distributions for both community categories nearly reached their 100%. Overall, we found that AI-claiming comments tend to emerge rapidly, and particularly in AI-disallowed communities, where they appear even more swiftly.
@@
Figure 3: The cumulative number of posts (y-axis) for time spanned between the publication of the first comment on a Suspected AI post and its first AI-claiming comment (x-axis).
@@
The time dynamics show a different response to suspected images in the two types of communities. To understand the response better, we conducted a sentiment analysis of comments suggesting AI involvement in the Suspected AI posts. To label the sentiment of comments implying AI, we used a large language model as classifier, following a similar approach as described in the previous sections (Wei et al. 2022; Rao et al. 2024). We used the LLM classifier to label the comments as Negative, Neutral, or Positive, using a prompt that can be found in Appendix D. To assess the accuracy of this method, we manually labeled a random sample of 100 comments and achieved an accuracy rate of 92% for this sentiment classification task.
@@ Discussion
Our study provides insights about how AI-generated content is impacting online communities, specifically those that deal with creative visual content—in this case, art-related subreddits.
@@
Perhaps surprisingly, we observed low volumes of (admitted, transparent) AI-generated content in these communities compared to other media-based posts. Across 2022 and 2023, years with popular discussion of AI and growing availability of image-generating tools, posts that either transparently used AI or were accused of doing so together amounted to less than 0.2% of media posts. Of course, it is possible that AI-generated image posts are more prevalent, but neither the authors nor the commenters are identifying them. Studies on the prevalence of AIGC online are sparse, but those that exist in other domains for now also indicate low numbers – for example, estimating 1.39% of machine-generated news articles for mainstream news websites (Hanley and Durumeric 2024) or up to 0.05% of AI-generated profiles on Twitter (Ricker et al. 2024). Of course, since this technology is still new, any presence of AIGC on public platforms points towards propagation in the future, and a fundamental shift to how people communicate digitally.
@@
Interestingly, though, our data shows that the volume of Transparent AI posts diminished over time after a first (expected rise), perhaps as novelty declined. On the other hand, responses to posts that are suspected of using AI rose later, and remained somewhat higher through the end of our data collection. These dynamics suggest that we have entered a new era in online communities, where the public is suspicious about the AI provenance of information. Unfortunately, we know that suspicion of AI content is likely to lower trust online more broadly (Jakesch, Hancock, and Naaman 2023). We also find that people may be becoming more negative about AIGC over time, as negative sentiments in comments for Suspected AI posts increased between early and late 2023. Additionally, our analysis highlights how this broad suspicion may interact with online community rules and moderation: AI-disallowed communities receive more suspicion comments, and these appear more quickly.
@@
In general, these findings raise the question of the role of rules and norms in online communities. There is clearly a difference in how communities decided to tackle the challenge of AI-generated content, and these norms and rules have both changed over time. It is important to continue to understand what brings about change of rules, how they are implemented, and how norms might shift as a result. Prior work identified that rules forbidding certain actions may appear later in a subreddit’s lifecycle and are often associated with larger subreddits (Reddy and Chandrasekharan 2023; Fiesler et al. 2018). It is likely that this trend will be the same for rules about AI-generated content. While we used a binary classification system to differentiate between AI-disallowed and AI-neutral communities, it is possible and even likely that our AI-disallowed communities implemented rules forbidding AI at some point in the middle of our dataset timespan. While this possibility makes the temporal trends more salient, it may underrepresent the strength of some of our comparative findings where, for example AI-disallowed communities have not actually disallowed AI for the full period. We see a rich space for future work here. For example, scholars could study communities immediately before and after a rule change forbidding AI is implemented. Future work may also look to evaluate the effects of AI how subreddit bans are implemented and enforced. For example, prior work in other setting has found that exposing rulesets more readily can decrease harmful content while increasing newcomer participation (Matias 2019).
@@
Our findings carry implications for how AI may enable newcomers to participate in online communities, but also how AI-interested newcomers may change existing communities. First, newcomers to communities create proportionally more Transparent AI posts than general image-based posts. This newcomer over-representation confirms prior work that suggests that the introduction of generative AI leads to more participation and posting (Guo et al. 2024; Wei and Tyson 2024; Zhou and Lee 2024). We show that users whose first image-based post to a subreddit is a self-labelled AI-generated image continue to post such AIGC at high rates within the subreddit. This behavior suggests that creators who initially joined these subreddits to transparently share AI-generated art, and have returned to post again, are truly interested in AI-generated images—or find it as their only possible avenue to contribute.  Kraut et al. (2012) identify five basic challenges that online communities face when dealing with newcomers, and our findings suggest that while the ready availability of AIGC may help with the recruitment and retention of new members, these new members do not become socialized in the exact same way as members that do not post AIGC. Instead, new members that post AIGC appear to wish to continue posting AIGC and may become turned off from online communities once AI content is banned. Indeed, we find evidence that when such newcomers post to subreddits where AI-generated content is prohibited, they are less likely to remain engaged within the community. Possible explanations for this lower retention include criticism or bans from the moderators of the subreddits, negative feedback from other participants, low interaction or engagement with their posts, or the possibility that these users are low-value contributors that are less invested in the community itself and thus contribute less frequently. Of course, lower retention of newcomers may not be a universal negative if new contributions are low-effort, and prior work has found moderators are sometimes willing to make these sacrifices to uphold community values (Lloyd, Reagle, and Naaman 2023).
@@
Our work also contributes an analysis of comments that suspect the author used AI to generate posts. The data suggests that high levels of community engagement on the post (e.g. comments) is associated with AI suspicion by commenters. In other words, AI suspicion may be more readily attached to successful posts. Our findings also suggest that AI accusations arrive rapidly, especially in AI-disallowed communities, reflecting the heightened awareness among participants regarding the possibility of generative art. This rapid response could indicate the role of moderators or community members working to uphold the standards and environment of their respective communities by quickly identifying and challenging suspected AI-generated works.
@@
This work also comes with some potential limitations. Of course, as we only looked at art communities, our findings mostly apply to this specific type of community and may not be broadly applicable. For example, perceptions, response and use of AIGC are likely to be different in non-visual or non-creative settings. Additionally, while we looked at 57 different subreddits, it is possible that our findings were disproportionately influenced by the most dominant communities—although we did our best to verify that this imbalance did not sway the results. Additionally, only data that was included in the PushShift torrents is included in this dataset, which overrepresents highly rated subreddits. Finally, the constraints of dataset collection means that it is possible that posts may have been deleted before they could be gathered by the torrents, though evidence suggests that PushShift tends to retrieve data quickly (Reagle 2023).
@@
Our findings correspond to a period of rapid advancements in AI-generated art technologies, which increasingly are able to produce high-quality content resembling traditional art forms. With this change, moderators of online communities contend with old challenges, such as whether appealing to newcomers is worth compromising on content quality, but are also faced with decisions that may fundamentally compromise their community values. Regardless of which rule sets communities choose to adopt, our findings show that suspicion and negativity over AI-generated content is here to stay. As the prevalence of AI-generated content grows, we must find ways to preserve and center core community values—such as creativity, authenticity, and openness—in social online ecosystems.
Appendix A Appendix A: List of Subreddits
@@
Appendix B Appendix B: List of Keywords
@@
Appendix C Appendix C: LLM Prompts for Data Classification
Title or Body of the Post
@@ Appendix C Appendix C: LLM Prompts for Data Classification
Comments Made by the Author of the Post
@@
Comments Made by Other Commenters
@@
Appendix D Appendix D: LLM Prompt for Sentiment Analysis

### v2

@@ 6.3. Feedback, Interactions and Sentiment
Figure 2. The cumulative distribution of posts across different upvote ratio values (x-axis), and the corresponding cumulative percentage of posts (y-axis) for different kinds of post (left vs right) and communities (top vs bottom).
@@
Given this observation, we further focused on the dynamics around Suspected AI posts, posts where a non-author commenter claimed or implied that the image attached to the post was generated using AI. To better understand how do community members react to posts containing media suspected of being AI-generated, we looked at the speed with which these claims emerged and the content of these claims. We began by calculating the time elapsed from the publication of the post until its first AI-claiming comment. Our analysis revealed that approximately 20% of the Suspected AI posts had received their first AI-claiming comment within one hour from posting. However, to compare between posts in the different types of communities, we had to control for the popularity of the post or its likelihood to be seen. To do that, we used the time window between the first comment on each Suspected AI post, and the first AI-claiming comment on that post. This between-comments data is presented in Figure 3, which shows the time between the comments (X-axis), and the cumulative number of posts (Y-axis). For instance, the leftmost data point in the figure indicates that for both community types, approximately 20% of the AI-claiming comments were made as the first comments on the posts (the X-axis tick mark of “1 minute” denoting “within one minute”, and mostly captured the first comment). The figure shows that these comments appear more quickly in AI-disallowed communities. For example, around 40% of the AI-claiming comments in these communities (red line) were published within one hour of the first comment. In AI-neutral communities, this speed of action occurred for less than 30% of the posts. The figure also shows that after approximately 36 hours, the distributions for both community categories nearly reached their 100%. Overall, we found that AI-claiming comments tend to emerge rapidly, and particularly in AI-disallowed communities, where they appear even more swiftly.
@@
Figure 3. The cumulative number of posts (y-axis) for time spanned between the publication of the first comment on a Suspected AI post and its first AI-claiming comment (x-axis).
@@
The time dynamics show a different response to suspected images in the two types of communities. To understand the response better, we conducted a sentiment analysis of comments suggesting AI involvement in the Suspected AI posts. To label the sentiment of comments implying AI, we used a large language model as a classifier, following a similar approach to that described in the previous sections, which is based on prior studies (Wei et al., 2022; Rao et al., 2024). We used the LLM classifier to label the comments as Negative, Neutral, or Positive, using a prompt that can be found in Appendix D, along with examples of input texts and their corresponding classifications. To assess the accuracy of this method, we manually labeled a random sample of 100 comments and achieved an accuracy rate of 92% for this sentiment classification task.
@@ 7. Discussion
Our study provides insights about how AI-generated content is impacting online communities, specifically those that deal with creative visual content – in this case, art-related subreddits.
@@
Perhaps surprisingly, we observed low volumes of (admitted, transparent) AI-generated content in these communities compared to other media-based posts (RQ1). Posts that either transparently used AI or were accused of doing so, together amounted to less than 0.5% of media posts. Of course, it is likely that AI-generated image posts are more prevalent, but neither the authors nor the commenters are identifying them. Studies on the prevalence of AIGC online are sparse, but those that exist in other domains for now also indicate low numbers – for example, estimating 1.39% of machine-generated news articles for mainstream news websites (Hanley and Durumeric, 2024) or up to 0.05% of AI-generated profiles on Twitter (Ricker et al., 2024). Of course, since this technology is still new, any presence of AIGC on public platforms points towards propagation in the future, and a fundamental shift to how people communicate digitally.
@@
Interestingly, though, our data reveals that the volume of Transparent AI posts (RQ1) decreased over time after a first (and expected) rise, perhaps due to declining novelty, changes in community policies and values, or shifts in how users present the content and respond to it. Most significantly, this decrease could be due to posters being more reluctant to admit their posts are AI-generated, as community rules and community and social norms around AI use are changing (Lloyd et al., 2025a). Indeed, responses to posts that are suspected of using AI rose later, and remained somewhat higher through the end of our data collection. These dynamics suggest that we have entered a new era in online communities, where the public is suspicious about the AI provenance of information. Unfortunately, we know that suspicion of AI content is likely to lower trust online more broadly (Jakesch et al., 2023). We also find that people may be becoming more negative about AIGC over time, as negative sentiments in comments for Suspected AI posts increased between early and late 2023.
@@
In general, these findings point to the dynamic nature of rules and norms in online communities. There is clearly a difference in how communities decided to tackle the challenge of AI-generated content, and research shows that these norms and rules about AI have both changed over time (Lloyd et al., 2025a). It is important to continue to understand what brings about change of rules, how they are implemented, and how norms might shift as a result. Prior work identified that rules forbidding certain actions may appear later in a subreddit’s lifecycle and are often associated with larger subreddits (Reddy and Chandrasekharan, 2023; Fiesler et al., 2018), and that this trend is similar for rules about AI-generated content (Lloyd et al., 2025a). While we used a binary classification system to differentiate between AI-disallowed and AI-neutral communities, it is possible and even likely that our AI-disallowed communities implemented rules forbidding AI at some point in the middle of our dataset timespan. This possibility may underrepresent the strength of some of our comparative findings where, for example, AI-disallowed communities have not actually disallowed AI for the full period. There is a rich space for future work in this area of shifting rules and norms. For example, scholars could study communities immediately before and after a rule change forbidding AI is implemented, and investigate the impact on content posting. Future work may also look to evaluate the effects of how subreddit bans are implemented and enforced. For example, prior work in other settings has found that exposing rulesets more readily can decrease harmful content while increasing newcomer participation (Matias, 2019).
@@
Our findings carry implications for how AI may enable newcomers to participate in online communities, but also how AI-interested newcomers may change existing communities (RQ2). First, we find that newcomers to communities create proportionally more Transparent AI posts than general image-based posts. This newcomer over-representation confirms prior work that suggests that the introduction of generative AI leads to more participation and posting (Guo et al., 2024; Wei and Tyson, 2024; Zhou and Lee, 2024). We show that users whose first image-based post to a subreddit is a self-labeled AI-generated image continue to post such AIGC at high rates within the subreddit. This behavior suggests that creators who initially joined these subreddits to transparently share AI-generated art, and have returned to post again, are truly interested in AI-generated images—or find it as their main possible avenue to contribute. Kraut et al. (Kraut et al., 2012) identify five basic challenges that online communities face when dealing with newcomers, and our findings suggest that while the ready availability of AIGC may help with the recruitment and retention of new members, these new members do not become socialized in the exact same way as members that do not post AIGC. Instead, new members that post AIGC appear to wish to continue posting AIGC and may become turned off from online communities once AI content is banned. Indeed, we find evidence that when such newcomers post to subreddits where AI-generated content is prohibited, they are less likely to remain engaged within the community. Possible explanations for this lower retention include criticism or bans from the moderators of the subreddits, negative feedback from other participants, low interaction or engagement with their posts, or the possibility that these users are low-value contributors that are less invested in the community itself and thus contribute less frequently. Of course, lower retention of newcomers may not be a universal negative if new contributions are low-effort, and prior work has found moderators are sometimes willing to make these sacrifices to uphold community values (Lloyd et al., 2025b).
@@
Our analysis of the response to content (RQ3) suggests that high levels of community engagement on the post (e.g. comments) are associated with AI suspicion by commenters, in what we may want to call Botwin’s Law (with apologies to Mr. Godwin): the more discussion or exposure an item gets, the more likely it is to be suspected of using AI. Or, in other words, AI suspicion may be more readily attached to successful posts. Our findings also suggest that AI accusations arrive rapidly, especially in AI-disallowed communities, reflecting the heightened awareness among participants regarding the possibility of generative art. This rapid response could indicate the role of moderators or community members working to uphold the standards and environment of their respective communities by quickly identifying and challenging suspected AI-generated works.
@@
We note that this work comes with some potential limitations. Of course, as we only looked at art communities, our findings mostly apply to this specific type of community and may not be broadly applicable. For example, perceptions, responses, and use of AIGC are likely to be different in non-visual or non-creative settings. Additionally, while we looked at 57 different subreddits, it is possible that our findings were disproportionately influenced by the most dominant communities—although we did our best to verify that this imbalance did not sway the results. For example, we verified that the observed trends in the timelines were not disproportionately influenced by the larger subreddits. However, potential biases toward more prominent or active communities could affect our conclusions, particularly if these dominant subreddits exhibit unique community rules, cultural norms, or dynamics. Additionally, only data that was included in the PushShift torrents is included in this dataset, which overrepresents highly popular subreddits. Finally, the constraints of dataset collection mean that it is possible that posts may have been deleted before they could be gathered by the torrents, though evidence suggests that PushShift tends to retrieve data quickly (Reagle, 2023).
The data processing and analysis steps may also have certain limitations, which could be addressed in future research. Identifying posts containing AI-generated media relies on text-based methods, which sometimes may capture user opinion and reactions, rather than accurately reflect the actual prevalence of AI-generated content. For example, in AI-neutral subreddits where AI-generated content is permitted, users can be less inclined to explicitly mention it in posts or comments. In contrast, in AI-disallowed subreddits, users may be more likely to claim that others’ content is AI-generated as it violates community rules. In the categorization of the communities step, given that AI-neutral communities do not explicitly address AI usage in their guidelines or descriptions, while AI-disallowed communities explicitly prohibit it, the observed differences in trends between these groups may reflect broader factors, such as more adaptive approach to emerging technologies by the AI-disallowed communities. That is, factors beyond the subreddit-specific policies regarding AI-generated content. Additionally, the analysis may face potential limitations related to the identification of newcomers, as the dataset may include bot or troll accounts considered as newcomers. Bots, in particular, are more likely to use AI-generated content (Grimme et al., 2022; Ferrara, 2023), and may be at higher risk of suspension by the platform. During data collection, measures were taken to filter out potential bot accounts based on usernames and repeated comments. However, these efforts may not have entirely excluded all accounts exhibiting bot-like behavior.
@@
In summary, our findings correspond to a period of rapid advancements in AI-generated art technologies, which increasingly are able to produce high-quality content resembling traditional art forms. With this change, moderators of online communities contend with old challenges, such as whether appealing to newcomers is worth compromising on content quality, but are also faced with decisions that may fundamentally compromise their community values. Regardless of which rule sets communities choose to adopt, our findings show that suspicion and negativity over AI-generated content is here to stay. As the prevalence of AI-generated content grows, we must find ways to preserve and center core community values – such as creativity, authenticity, and openness – in social online ecosystems.
Appendix A List of Subreddits
@@
Appendix B List of Keywords
@@
Appendix C LLM Prompts for Data Classification
C.1. Title or Body of the Post
@@
C.2. Comments Made by the Author of the Post
@@
C.3. Comments Made by Other Commenters
@@
Appendix D LLM Prompt for Sentiment Analysis